# CHASE: A Causal Heterogeneous Graph based Framework for Root Cause Analysis in Multimodal Microservice Systems

### v1

@@ Abstract
In recent years, the widespread adoption of distributed microservice architectures within the industry has significantly increased the demand for enhanced system availability and robustness. Due to the complex service invocation paths and dependencies at enterprise-level microservice systems, it is challenging to locate the anomalies promptly during service invocations, thus causing intractable issues for normal system operations and maintenance. In this paper, we propose a Causal Heterogeneous grAph baSed framEwork for root cause analysis, namely CHASE, for microservice systems with multimodal data, including traces, logs, and system monitoring metrics. Specifically, related information is encoded into representative embeddings and further modeled by a multimodal invocation graph. Following that, anomaly detection is performed on each instance node with attentive heterogeneous message passing from its adjacent metric and log nodes. Finally, CHASE learns from the constructed hypergraph with hyperedges representing the flow of causality and performs root cause localization. We evaluate the proposed framework on two public microservice datasets with distinct attributes and compare with the state-of-the-art methods. The results show that CHASE achieves the average performance gain up to 36.2%(A@1) and 29.4%(Percentage@1), respectively to its best counterpart.
@@ I Introduction
Microservice architecture has recently surged in popularity due to its significant advantages for contemporary industrial service-oriented systems. This architectural design involves a collection of small, independent, and loosely coupled applications, where each application delivers a specific business function [1]. Specifically, each microservice is independently developed, deployed, and managed by a dedicated team, with communication facilitated through APIs and protocols such as REST, HTTP, and messaging systems. It is widely acknowledged that microservice architecture enhances flexibility, scalability, and fault tolerance [2, 3, 4].
However, fault localization and analysis in microservice systems often pose significant challenges due to high complexity [5]. In reality, operation engineers and developers are frequently required to utilize data such as distributed invocation histories and logging tools to quickly identify and isolate issues. In complex systems where microservices are interconnected, each method or function call might trigger multiple distributed service invocations, either synchronously or asynchronously. This complexity can overwhelm operation engineers, especially when they need to manually gather data from various sources during system failures, thereby complicating the issue identification and resolution process. Implementing trace, metric, and log functionalities, and understanding how to use these data collectively, are essential for effective management of complex microservice systems. Tracing allows for the tracking of a request’s path across various microservices, offering valuable insights into performance and identification of bottlenecks [2]. Metrics, sometime being refered as system monitoring metrics, provide quantitative data related to system performance, such as response times and error rates, which are crucial for performance optimization and capacity planning [6]. System logs, on the other hand, capture detailed information about system events and transactions, facilitating debugging and troubleshooting processes [7]. Together, traces, metrics, and logs create a comprehensive monitoring system that encompasses all necessary multimodal information, which helps enhance visibility into the system’s behavior, thereby supporting performance monitoring, debugging, and ensuring overall system reliability.
Fig. 1 illustrates the execution process of service invocations within an enterprise-level microservice system, with trace, metrics, and logs data being monitored and recorded. The microservice trace data is visualized as a directed graph where each node represents an instance of various services such as network gateways, message queues, databases, and other business-specific services. Directed edges show the flow of requests between services and application instance. Metric data, such as response times or error rates, are captured as time-series data, illustrating changes over time. Logs are typically formatted according to the configuration of the log system and provide textual insights into instance activities, which can be analyzed using natural language processing techniques to extract semantic information.
As the complexity of systems continues to grow, there is an increasing emphasis on integrating Root Cause Analysis (RCA) methods to timely uncover the underlying reasons behind issues or failures [8, 9, 10]. RCA is particularly crucial in microservice systems due to their intricate and dynamic interactions between various service instances. When failures occur in large-scale microservice systems, they can spread rapidly, leading to significant disruptions, as evidenced by the 12-hour outage experienced by Didi, a major Chinese ride-hailing platform, in November 2023, resulting in a loss of 56.43 million [11]. RCA in microservice systems involves not just identifying general error types for further investigation but also pinpointing specific instance nodes along with their problematic logs and metrics [12, 13]. For instance, as illustrated in Fig. 1, a typical RCA framework collects the runtime information from the microservice system and outputs its error diagnosis, indicating the DB instance node circled in red as the root cause of the system failure. Additionally, the RCA framework outputs several logs from the error stack and detected anomaly metrics to facilitate further manual checks. However, traditional RCA methods only offer suggestive clues, necessitating detailed manual analysis to determine the exact root cause [14]. Furthermore, with the ongoing evolution of monolithic microservice architectures, new challenges that require handling multimodal information at runtime and the demands of robust anomaly diagnosis arise, underscoring the critical importance of effective RCA as a pressing field.
@@
To date, numerous models have been proposed for root cause analysis. One class of approaches explores the feasibility of building causality graphs using merely trace information, which consists of service invocation chains from various service instances. Based on such intuition, causal analysis algorithms such as PC [15] and GES [16] are employed to infer the causal relationships between service instances in the graph. Following that, a line of research [17, 18, 19, 20, 21, 22] starts to integrate root cause analysis with machine learning models and graph neural networks (GNNs), from which the graph learning techniques are leveraged to model the pairwise bivariate correlation of instances connected with directed invocation edges. In these work, multimodal data of service instances are used as supplementary features to enrich features such as node embeddings and edge weights, while the uncovering of bivariate correlation between pairs of instance nodes are treated as the homogeneous graph learning task in the trace typology, leading to the following challenges:
@@
    Information heterogeneity of microservice systems. The microservice system information composed of log, metric and trace should be considered as multimodal data. The log information contains systemically defined natural languages. The metric information collects time series data of numerical values. And the trace information is the typology of the invocation graph. With multimodal information involved, it is both reasonable and natural to explore the heterogeneity attributes of the service invocation graph.
@@
    Multivariate causal correlation between instances. Since each instance node affects all its downstream invocation, the causal information of such instance node can possibly propagate to nodes that lies multiple hops away from itself. It is challenging to capture the causality flow using graph convolution networks, which primarily focus on modeling pairwise locality correlation of directed edges. Instead, more sophisticated graph structures that can model multivariate correlations among subsets of nodes with varying cardinality should be considered.
To tackle these issues, we propose a novel causal heterogeneous graph based framework named CHASE, which is applicable to inductively learn the root cause anomaly of microservice systems with the presence of trace, log, and metric data all together. Based on multimodal data, a heterogeneous invocation graph is constructed using trace topology, with additional metric and log nodes connected to the source instance nodes where the information is gathered. Representative node embeddings are then generated using designated encoders for each multimodal data type. The procedure for instance-level anomaly detection is carried out using heterogeneous attentive message passing. By doing so, CHASE can accurately locate root cause instances with excellent performance by learning from the constructed hypergraph, with hyperedges representing causality propagation in the trace. We evaluate CHASE on two datasets with different attributes—one with static trace topology and one with dynamic trace topology. Our experimental results show that CHASE significantly improves fault localization accuracy compared to several state-of-the-art methods.
@@
    We propose a novel graph learning framework named CHASE to handle multimodal data of log, metric and trace under the task of root cause analysis, in which encode trace, metrics and log information are encoded along with the construction of invocation graph-based typology. We accomplish multimodal feature fusion by solving the instance-level anomaly detection with heterogeneous message passing. Hereby, we attain the modeling of multimodal data and information heterogeneity of microservice systems.
@@
    We conduct extensive experiments using two public datasets from microservice systems and compare the performance with a number of traditional and GNN-based baselines, demonstrating that our proposal can outperform comparative methods in the root cause analysis task.
@@ II-B GNN based Root Cause Analysis
Recent advancement of GNNs has unleashed great potential in analyzing the root cause problems, especially for cases relying on graph structure data. Owing to the natural compatibility with generated causal graphs, researchers started to leverage GNN to learn the failure propagation patterns in the graph, thus giving a stronger generalization ability compared with random walk based RCA frameworks on causal graphs generated with Peter-Clark(PC) algoritm [22]. CausalRCA [21] designs a gradient-based causal structure learning to capture linear and non-linear causal relations in monitoring metrics and outputs a weighted causal directed acyclic graph (DAG). Diagfusion [22] reports the state-of-the-art RCA results using GNN. It combines deployment data and traces to build a dependency graph and uses GNN to generate the embeddings for each service instance, which are then used to achieve two-fold failure diagnosis, i.e., root cause fault localization and failure type determination. A hierarchical causal network framework named REASON[31] is applied to model the fault propagation of the microservice trace with both within-network and across-network causal relationships for root cause localization. Random walk on the learnt causal network is further applied to locate a system fault. It is worth noting that CausalRCA uses the encoder-decoder structure for unsupervised learning to derive the weight of edges, which is considered insufficient to represent the significance between causes and effects in the causal graph. Diagfusion, on the other hand, applies graph convolutional network for representation learning on invocation graphs in an end-to-end fashion, requiring a certain amount of labelled data.
@@ III CHASE Framework
In this section, we elaborate the CHASE framework with three stages, including invocation graph generation, instance anomaly detection with heterogeneous message passing and root cause analysis with hypergraph learning. Fig. 2 depicts the overall architecture. We list out all necessary notations used in this paper in Table I.
@@ III-A Multimodal Invocation Graph Construction
In a microservice system, as shown in Fig. 1, the process of handling a request is usually reflected with a microservice trace, which can be modeled by a DAG with instances as nodes and each directed edge $e(u,v)$ representing that instance $v$ is called by instance $u$. There are multiple types of service instances, including application instances, message instances, gateway instances, database instances. The instances and invocation edges on a trace are extracted and assembled to obtain the complete microservice invocation graph to represent business processes. In general, each instance will continuously output logs during runtime, while metrics of the instance’s container and physical server, will also be recorded by the DevOps platform simultaneously.
@@
Note that the raw feature of log nodes are natural language texts with semantic meaning, features of metric nodes are composed of real valued time series data and the instance nodes have categorical features. As a result, invocation graph with log, metric and instance nodes can be considered as a graph with multimodal information. Hence, we apply respective encoders to encode the feature of log, metric and instance nodes into embeddings denoted as $\mathbb{E}(\mathcal{L}),\mathbb{E}(\mathcal{M}),\mathbb{E}(\mathcal{I})$, respectively:
@@ III-B Instance Level Anomaly Detection with Heterogeneous Message Passing
After the whole encoding process where all the log node embeddings, metric node embeddings and instance node embeddings are encoded into $\mathbb{E}(\mathcal{L})$, $\mathbb{E}(\mathcal{M})$, $\mathbb{E}(\mathcal{I})$, we construct the multimodal invocation graph. The task of instance level anomaly detection is a conventional procedure in root cause analysis. In this section, we handle this task based on the intrinsic heterogeneity of the invocation graph, which is inspired by the architecture design of heterogeneous graph transformer.
@@ III-B1 Attentive Anomaly Score
For each instance node in the microservice system, multiple types of runtime metrics and logs are continuously produced and monitored by the DevOps platform. It is crucial to identify which of these monitored records are the most representative for the occurrence of the detected instance-level anomaly. We quantify this anomaly score, denoted as $\mathcal{AS}$, with the heterogeneous mutual attention weight between the instance node and all its neighbors in the invocation graph.
@@
Both $A_{\mathcal{L}-\mathcal{I}}$, $A_{\mathcal{M}_{n}-\mathcal{I}}$ are of dimension $d/H\times d/H$, which represents the projection matrix for calculating the attention weight between the instance node and its multimodal neighbors. This matrix is then to be shareable among all attention heads. $\phi_{\mathcal{L}}$ and $\phi_{\mathcal{M}_{n}}$ are the prior significance granted to different types of metrics and logs, which can be learnable and initialized to ones, inferring our prior assumption that all kinds of logs and metrics are equally important to the instance-level anomaly.
@@ III-B3 Instance Level Anomaly Embedding Update
III-C Root Cause Analysis with Hypergraph Learning
@@ IV-A1 DataSets
We give the details of the datasets here. The first dataset is a public dataset named Generic AIOps Atlas, provided by Cloudwise [^2] https://github.com/CloudWise-OpenSource/GAIA-DataSet; the second dataset is a real-world dataset collected from the AIOps 2020 competition [^3] https://github.com/NetManAIOps/AIOps-Challenge-2020-Data.
@@ IV-A3 Evaluation Metrics
TABLE II: Overall Performance Comparison
DataSet Evaluation Metric Causality-based Correlation-based GNN-based
PC GES CloudRanger MicroRCA TrinityRCL CausalRCA DiagFusion Proposed Method
GAIA A@1 0.2960 0.3003 0.3290 0.3421 0.4503 0.3652 0.4121 0.6135
A@3 0.6368 0.5399 0.4771 0.5528 0.8244 0.4973 0.8157 0.8823
Avg@5 0.5953 0.5154 0.4883 0.5712 0.7651 0.5966 0.7478 0.8276
AIOps 2020 Percentage@5 - 0.11 - 0.08 0.12 - 0.12 0.15
Percentage@3 - 0.08 - 0.13 0.14 - 0.14 0.16
Percentage@1 - 0.14 - 0.17 0.15 - 0.14 0.22
@@ IV-D Ablation Study
Considering the V2 variant of CHASE, we replace the heterogeneous message passing layer with a homogeneous graph attention layer, which results in a 37% performance drop in A@$1$. This emphasizes the significance to model the heterogeneous correlation of the invocation graph. To provide a clearer illustration of the impact of learning the heterogeneity attributes among metrics, logs, and instances through message passing in CHASE, Fig. 4 plots a comparison between the homogeneous edge weight and the heterogeneous edge weight on the $160^{th}$ trace in GAIA dataset (webservice2 as the root cause instance). As the edges related to webservice2 (last columns in both figures) are granted with higher weights, it indicates that heterogeneous message passing layer is able to enhance the root cause inference’s performance, which is achieved by directing more attention towards the edges related to the instance of the root cause.
In the context of causality inference, it has been observed that directly adding the detection head to the output of the heterogeneous message passing layer in the V3 variant causes the causal hypergraph to degrade into a heterogeneous graph neural network. In this scenario, information propagates only through K hops (where K represents the number of layers), which contradicts the nature of anomaly causality that may propagate across an arbitrary number of hops depending on the invocation. This leads to a performance drop of 23.7% in A@@@@1.

### v2

# CHASE: A Causal Hypergraph based Framework for Root Cause Analysis in Multimodal Microservice Systems

@@ Abstract
In recent years, the widespread adoption of distributed microservice architectures across the industry has greatly increased the demand for improved system availability and robustness. Enterprise-level microservice systems often involve complex service invocation paths and system dependencies, making it challenging to quickly identify anomalies during service interactions, which inevitably hinders normal system operations and maintenance. In this paper, we propose CHASE—a Causal Hypergraph grAph baSed framEwork—for root cause analysis in microservice systems. CHASE is designed to handle multimodal data sources, including traces, logs, and system monitoring metrics. Specifically, CHASE begins by encoding each type of input data—such as traces, logs, and metrics—into representative embeddings. These embeddings are then integrated into a unified multimodal invocation graph that captures the relationships among services. Anomaly detection is subsequently applied to each service instance node through attentive heterogeneous message passing, leveraging contextual information from adjacent metric and log nodes. Finally, CHASE constructs a causal hypergraph, where hyperedges represent inferred causal relationships among anomalies, and utilizes this structure to accurately localize root causes within the microservice system. We evaluate the proposed framework on two public microservice datasets with distinct attributes and compare with seven benchmark methods. The results show that CHASE achieves the average performance gain up to 36.2%(A@1) and 29.4%(Percentage@1), respectively to its best counterpart.
@@ I Introduction
Microservice architecture has recently surged in popularity due to its significant advantages such as enhancing flexibility, scalability, and fault tolerance for contemporary industrial service-oriented systems [1, 2, 3]. This architectural design involves a collection of small, independent, and loosely coupled application instances, where each application delivers a specific business function [4]. Specifically, each microservice is independently developed, deployed, and managed by a dedicated team, with communication facilitated through APIs and protocols such as REST, HTTP, and messaging systems.
However, fault localization and analysis in microservice systems often pose significant challenges as microservices are interconnected and each method or function call might trigger multiple distributed service invocations, either synchronously or asynchronously [5, 6, 7]. Operations engineers are often overwhelmed by the need to manually collect data from diverse sources during system failures. This process involves managing intricately structured function call information, semi-structured log text, and system performance metrics, thereby complicating the identification and resolution of issues. As the complexity of systems continues to grow, there is an increasing emphasis on integrating Root Cause Analysis (RCA) methods to timely uncover the underlying reasons behind issues or failures [8, 9, 10].
@@
Traditional RCA methods only offer suggestive clues, necessitating detailed manual analysis to determine the exact root cause [11]. To date, numerous models have been proposed for root cause analysis. One class of approaches explores the feasibility of building causality graphs using merely trace information, which consists of service invocation chains from various service instances. Based on such intuition, causal analysis algorithms such as PC [12] and GES [13] are employed to infer the causal relationships between service instances in the graph. Following that, a line of research [14, 15, 16, 17, 18, 19, 20, 21, 6, 22, 23] starts to integrate root cause analysis with machine learning models and graph neural networks (GNNs), from which the graph learning techniques are leveraged to model the pairwise bivariate correlation of instances connected with directed invocation edges. In these work, multimodal data of service instances are used as supplementary features to enrich features such as node embeddings and edge weights, while the uncovering of bivariate correlation between pairs of instance nodes are treated as the homogeneous graph learning task in the trace typology, leading to the following challenges:
@@
    Joint processing of multimodal information and invoke topology. Existing non-graph neural network frameworks primarily extract causal relationships from multimodal data to construct causal structure graphs but fail to fully utilize call topology information. While GNN frameworks adopt graph-based learning methods, they fall short in simultaneously capturing multimodal information, instance invoking topology, and causal propagation.
@@
    Multivariate causal correlation capture among instances. In microservice systems, each instance node can influence multiple downstream calls, and causal information may propagate across nodes several hops away. Traditional graph neural networks primarily focus on modeling pairwise local correlations along directed edges, making it challenging to capture such complex causal flows.
To tackle these issues, we propose a novel causal hypergraph-based framework named CHASE, which is applicable to inductively learn the root cause anomaly of microservice systems with the presence of trace, log, and metric data all together. Based on multimodal data, a heterogeneous invocation graph is first constructed using trace topology, with additional metric and log nodes connected to the source instance nodes where the information is gathered. Representative node embeddings are then generated using designated encoders for each multimodal data type. The procedure for instance-level anomaly detection is carried out using heterogeneous attentive message passing. By doing so, CHASE can accurately locate root cause instances with excellent performance by learning from the constructed hypergraph, with hyperedges representing causality propagation in the trace. We evaluate CHASE on two datasets with different attributes—one with static trace topology and one with dynamic trace topology. Our experimental results show that CHASE significantly improves fault localization accuracy compared to several state-of-the-art methods.
@@
    We introduce CHASE, an innovative graph learning framework tailored to process multimodal data, including logs, metrics, and traces, specifically for root cause analysis. This framework integrates the encoding of multimodal information with the development of an invocation graph-based typology. We accomplish multimodal feature fusion by solving the instance-level anomaly detection with heterogeneous message passing. Hereby, we attain the modeling of multimodal data and information heterogeneity of microservice systems.
@@
    We conduct extensive experiments using two public datasets from microservice systems and compare the performance with a number of traditional and GNN-based baselines, demonstrating that our proposal can outperform comparative methods in the root cause analysis task. Apart from that, we release the source code [^1] https://drive.google.com/file/d/11erha3k8FeA67z-sfKGReqHz66PpO6o4/view?usp=sharing to advance the developement of this field.
@@ II-B GNN based Root Cause Analysis
Recent advancement of GNNs has unleashed great potential in analyzing the root cause problems, especially for cases relying on graph structure data. Owing to the natural compatibility with generated causal graphs, researchers started to leverage GNN to learn the failure propagation patterns in the graph, thus giving a stronger generalization ability compared with random walk based RCA frameworks on causal graphs generated with Peter-Clark(PC) algoritm [19]. CausalRCA [18] designs a gradient-based causal structure learning to capture linear and non-linear causal relations in monitoring metrics and outputs a weighted causal directed acyclic graph (DAG). Diagfusion [19] reports the state-of-the-art RCA results using GNN. It combines deployment data and traces to build a dependency graph and uses GNN to generate the embeddings for each service instance, which are then used to achieve two-fold failure diagnosis, i.e., root cause fault localization and failure type determination. A hierarchical causal network framework named REASON[34] is applied to model the fault propagation of the microservice trace with both within-network and across-network causal relationships for root cause localization. Random walk on the learnt causal network is further applied to locate a system fault. It is worth noting that CausalRCA uses the encoder-decoder structure for unsupervised learning to derive the weight of edges, which is considered insufficient to represent the significance between causes and effects in the causal graph. Diagfusion, on the other hand, applies graph convolutional network for representation learning on invocation graphs in an end-to-end fashion, requiring a certain amount of labelled data. Moreover, Eadro[35] employs GAT to learn representations of multimodal information status. It integrates anomaly detection and root cause localization into a unified end-to-end model, facilitating knowledge sharing and joint training.
@@ III CHASE Framework
In this section, we elaborate the CHASE framework with three stages, including invocation graph generation, instance anomaly detection with heterogeneous message passing, and root cause analysis with hypergraph learning. Fig. 2 depicts the overall architecture. Our proposed framework first constructs a graph model by integrating multimodal data and topological dependencies (Fig. 2(a)). Then, a heterogeneous graph transformer (HGT) is employed to perform representation learning for instance nodes and their associated data nodes (Fig. 2(b)), aiming to incorporate the heterogeneity attributes from multimodal data. Finally, hypergraph convolution is utilized to conduct root cause localization for the instance nodes (Fig. 2(c)). We list all necessary notations used in this paper in Table I and provide a more detailed explanation of the three submodules in the following subsections.
@@ III-A Multimodal Invocation Graph Construction
In a microservice system, as illustrated in Fig. 1, the handling of a request is typically represented by a microservice trace, which can be modeled as a directed acyclic graph (DAG). In this representation, instances serve as nodes, and each directed edge $e(u,v)$ signifies that instance $v$ is invoked by instance $u$. There are multiple types of service instances, including application instances, message instances, gateway instances, database instances. The instances and invocation edges on a trace are extracted and assembled to obtain the complete microservice invocation graph to represent business processes. Generally, each instance continuously generates logs during its runtime, while the DevOps platform concurrently records metrics related to the instance’s container and physical server.
@@
Note that the raw feature of log nodes are natural language texts with semantic meaning, features of metric nodes are composed of real valued time series data and the instance nodes have categorical features. Consequently, an invocation graph that includes log, metric, and instance nodes can be regarded as a graph containing multimodal information. Hence, we apply respective encoders to encode the feature of log, metric and instance nodes into embeddings denoted as $\mathbb{E}(\mathcal{L}),\mathbb{E}(\mathcal{M}),\mathbb{E}(\mathcal{I})$, respectively:
@@ III-B Instance Level Anomaly Detection with Heterogeneous Message Passing
After the whole encoding process where all the log node embeddings, metric node embeddings and instance node embeddings are encoded into $\mathbb{E}(\mathcal{L})$, $\mathbb{E}(\mathcal{M})$, $\mathbb{E}(\mathcal{I})$, we construct the multimodal invocation graph. The task of instance level anomaly detection is a conventional procedure in root cause analysis. In this section, we handle this task based on the intrinsic heterogeneity of the invocation graph inspired by the architecture design of heterogeneous graph transformer.
@@ III-B1 Attentive Anomaly Score
For each instance node in the microservice system, multiple types of runtime metrics and logs are continuously produced and monitored by the DevOps platform. It is crucial to identify which of these monitored records are most representative of the detected instance-level anomaly. We achieve this by quantifying the anomaly score, denoted as $\mathcal{AS}$, with the heterogeneous mutual attention weight between the instance node and all its neighbors in the invocation graph.
@@
Both $A_{\mathcal{L}-\mathcal{I}}$, $A_{\mathcal{M}_{n}-\mathcal{I}}$ are of dimension $d/H\times d/H$, which represents the projection matrix for calculating the attention weight between the instance node and its multimodal neighbors. This matrix is designed to be shared among all attention heads. The parameters $\phi_{\mathcal{L}}$ and $\phi_{\mathcal{M}_{n}}$ represent the prior significance assigned to different types of metrics and logs. These parameters are learnable and can be initially set to ones, reflecting our initial assumption that all types of logs and metrics are equally important for identifying instance-level anomalies.
@@ III-B3 Instance Level Anomaly Embedding Update
III-C Hypergraph Causality Learning for RCA
@@ IV-A1 DataSets
The first dataset is a public dataset named Generic AIOps Atlas, provided by Cloudwise [^3] https://github.com/CloudWise-OpenSource/GAIA-DataSet; the second dataset is a real-world dataset collected from the AIOps 2020 competition [^4] https://github.com/NetManAIOps/AIOps-Challenge-2020-Data. We give the details of the datasets here.
@@ IV-A2 Baseline Methods
TABLE II: Overall Performance Comparison
Dataset GAIA AIOps 2020
Evaluation Metric A@1 A@3 Avg@5 Percentage@5 Percentage@3 Percentage@1
Causality-based PC 0.2960 0.6368 0.5953 - - -
GSE 0.3003 0.5399 0.5154 0.11 0.08 0.14
Correlation-based CloudRanger 0.3290 0.4771 0.4883 - - -
MicroRCA 0.3421 0.5528 0.5712 0.08 0.13 0.17
TrinityRCL 0.4503 0.8244 0.7651 0.12 0.14 0.15
GNN-based CausalRCA 0.3652 0.4973 0.5966 - - -
DiagFusion 0.4121 0.8157 0.7484 0.12 0.14 0.14
Proposed Method 0.6135 0.8823 0.8276 0.15 0.16 0.22
@@ IV-D Ablation Study
Considering the V2 variant of CHASE, we replace the heterogeneous message passing layer with a homogeneous graph attention layer, which results in a 37% performance drop in A@$1$. This emphasizes the significance to model the heterogeneous correlation of the invocation graph. To provide a clearer illustration of the impact of learning the heterogeneity attributes among metrics, logs, and instances through message passing in CHASE, Fig. 4 plots a comparison between the homogeneous edge weight and the heterogeneous edge weight on the $160^{th}$ trace in GAIA dataset (webservice2 as the root cause instance). As the edges related to webservice2 (last columns in both figures) are granted with higher weights, it indicates that the heterogeneous message passing layer is able to enhance the root cause inference’s performance, which is achieved by directing more attention towards the edges related to the instance of the root cause.
In the context of causality inference, it has been observed that directly integrating the detection head into the output of the heterogeneous message passing layer in the V3 variant leads to the causal hypergraph degrading into a heterogeneous graph neural network. In this scenario, information propagates only through K hops (where K represents the number of layers), which contradicts the nature of anomaly causality that may propagate across an arbitrary number of hops depending on the invocation. This leads to a performance drop of 23.7% in A@@@@1.